# Extremal correlation coefficient for functional data

### v2

@@ 5 A simulation study
Assume that curves are observed on the regularly spaced grid $\{j/J,\ j\in\{1,\ldots,J\}\}$ on $[0,1]$, with each point assigned equal weight. The inner product and norm in $L^{2}([0,1])$ are then given by
@@
$g_{X}(x)=\frac{x}{\|x\|^{1-{\alpha_{X}}/\alpha}},\ \ \ \ g_{Y}(y)=\frac{y}{\|y\|^{1-{\alpha_{Y}}/\alpha}},\ \ \ \ x,y\in L^{2},$ (16)
where $\alpha$ is a desired tail index. Applying $g_{X}$ and $g_{Y}$ to $X$ and $Y$, respectively, ensures that $P(\|g_{X}(X)\|>\cdot)$ and $P(\|g_{Y}(Y)\|>\cdot)$ are regularly varying with the same index $-\alpha$. Since this method adjusts only the scale of curves, the transformed curves still retain their original shapes.
Step 4. Given tail equivalent marginals, take the $k$ largest pairs of $[X_{i},Y_{i}]^{\top}$, based on their norm, i.e., $R_{i}=\|X_{i}\|\vee\|Y_{i}\|$, and then compute $\hat{\rho}_{n,k}$ as in (14). By Lemma A.3 (i), if $[X,Y]^{\top}$ are regularly varying in $L^{2}\times L^{2}$ with index $-\alpha$, then $R=\|X\|\vee\|Y\|$ is regularly varying in $\mathbb{R}_{+}$ with the same index. Therefore, we choose $k$ that results in successful tail estimation for $R$ in finite samples. In the literature on tail estimation, various methods for selecting $k$ have been introduced, e.g., Hall & Welsh (1985)Hall (1990)Drees & Kaufmann (1998),  Danielsson et al. (2001), just to name a few. We use the method of  Danielsson et al. (2016), which is implemented using the function mindist of the R package tea.
We now outline the design of our simulation study. We generate functional observations in such a way that the theoretical value of $\rho_{XY}$ can be computed analytically, so that we can see how close $\hat{\rho}_{n,k}$ is to the true value. Suppose that $Z_{1}$ and $Z_{2}$ are i.i.d. random variables in ${\mathbb{R}}$ satisfying ${\rm pr}(|Z_{1}|>z)=z^{-\alpha}$ with equal chance of $Z_{1}$ being either negative or positive. Also, let $N_{1}$, $N_{2}$, and $N_{3}$ be i.i.d. normal random variables in ${\mathbb{R}}$ with mean 0 and variance 0.5. Consider $\{\phi_{j},j\geq 1\}$ defined by (3) and recall that it is an orthonormal basis in $L^{2}([0,1])$. These functions are simulated on a grid of 100 equally–spaced points on the unit interval $[0,1]$. We consider the following data generating processes, for $-1\leq\rho\leq 1$,
$X(t)=Z_{1}\phi_{1}(t)+N_{1}\phi_{2}(t)+N_{2}\phi_{3}(t);\ \ Y(t)=\rho Z_{1}\phi_{1}(t)+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}(t)+N_{3}\phi_{3}(t).$ (17)
@@
Let $[Z_{1},Z_{2}]^{\top}$ be a random vector in ${\mathbb{R}}^{2}$ consisting of i.i.d. components $Z_{j}$ whose magnitude is regularly varying with tail index $\alpha>2$, i.e., for some $c_{+}$, $c_{-}\geq 0$, ${\rm pr}(Z_{1}>z)\sim c_{+}z^{-\alpha}$ and ${\rm pr}(Z_{1}<-z)\sim c_{-}z^{-\alpha}$, where $f(z)\sim g(z)$ if and only if $\lim_{z\to\infty}f(z)/g(z)=1$. Also, let $\{\phi_{j},\ j\geq 1\}$ be a set of orthonormal elements in ${\mathbb{S}}$. Then, for $X$ and $Y$ in (17),
@@
We consider $\rho_{XY}\in\{0,\pm 0.1,\pm 0.2,\ldots,\pm 0.9,\pm 1\}$ and $\alpha\in\{3,4,5\}$, from which values of $\rho$ can be obtained by Lemma 5.1. For each $\rho$, we generate $[X_{i},Y_{i}]^{\top}$, $1\leq i\leq n$, that are i.i.d. copies of $[X,Y]^{\top}$, with sample sizes $n\in\{100,500,2000\}$. In each case, 1000 replications are generated. By Proposition 7.1 and Example 7.3 of Meinguet & Segers (2010), $X$ and $Y$ in (17) are regularly varying with the same index $-\alpha$, so we proceed directly to Step 4 and compute $\hat{\rho}_{n,k}$. When choosing $k$, we also consider an alternative approach based on the KS distance, as introduced by Clauset et al. (2009). The key difference from Danielsson et al. (2016) is that Clauset et al. (2009) computes the distance from tail distributions, rather than tail quantiles. The method is implemented using poweRlaw R package.
@@
We report the magnitude of empirical biases (the absolute difference between the average and the theoretical value), along with standard errors computed as the sample standard deviations. Using the optimal $k$s selected by the method from Danielsson et al. (2016), the results are shown in Table 2 for $\alpha=3$. The results for $\alpha\in\{4,5\}$, as well as those using the method from Clauset et al. (2009), are provided in Section D.1 of Supplementary material, as they lead to similar conclusions. The conclusions are summarized as follows.
1. The estimator is consistent as the biases decrease with increasing sample sizes, across nearly all values of $\rho_{XY}$ and for all $\alpha\in\{3,4,5\}$ considered. However, the finite performance using the optimal $k$s appears to depend on the tail index $\alpha$, with smaller $\alpha$ resulting in small biases.
2. The bias tends to increase in magnitude as $|\rho_{XY}|$ approaches 1. This could be due to the effect of the boundary, $\rho_{XY}\in\{-1,1\}$. These barriers cause the bias to increase, although it gets lower again at the boundary.
3. The standard errors are observed to be non-uniform across $\rho_{XY}$, they roughly behave like a quadratic function of $\rho_{XY}$ with its peak at ±.$\pm.5$.
@@ 6.1 Extremal dependence of intraday returns on sector ETFs
Ticker Sector $\hat{\alpha}$
XLY Consumer Discretionary 3.8
XLP Consumer Staples 2.6
XLE Energy 4.2
XLF Financials 4.0
XLV Health Care 3.9
XLI Industrials 3.7
XLB Materials 3.4
XLK Technology 4.7
XLU Utilities 3.8
@@
Figure 3 shows estimates of the pairwise extremal correlation coefficient across the nine ETF sectors. All pairs exhibit positive extremal correlations ($\hat{\rho}_{n,k}=0.43\sim 0.96$), and 44% of the pairs have strong extremal correlations above 0.7. We see that the CIDRs overall exhibit matching patterns of cumulative intraday returns on extreme market volatility days during the COVID-19 market turbulence, where most sectors either drop or increase together. However, our coefficient reveals more subtle information as well. For example, extreme return curves of XLF (Financials) are exceptionally strongly correlated with extreme curves for XLV, XLB and XLK (Health Care, Materials, Technology), but moderately correlated with XLU (Utilities). We do not aim at an analysis of the stock market or the economy, but we note that some findings are interesting. One might expect that the financial sector (mostly banks) to be strongly affected by the technology sector (mostly large IT companies like Google or Microsoft) because such mega corporations dominate the U.S. stock market. The similarity of extreme return curves for XLF and XLK is shown in the leftmost panels of Figure 2. One might expect bank stocks to be less affected by utility companies, whose revenues are largely fixed, but it’s less obvious why banks are strongly correlated with Health Care and Materials sectors. As another comparison, XLY (Consumer Discretionary) and XLU (Utilities) show a moderate extremal correlation of 0.43. Their extreme curves exhibit relatively dissimilar patterns, as seen in the middle panels in Figure 2.
@@
We conclude by emphasizing that our tool offers a more precise quantification of intraday risk during extreme events. First, by analyzing curve shapes, it provides a better assessment of intraday risks. The left plot of Figure 4 displays the pairwise coefficients from single-valued closing returns. This plot reveals somewhat different information from Figure 3. For instance, between XLY and XLE, the closing returns show a $-0.3$ correlation, indicating negative correlation during extreme days. However, this value does not accurately reflect the positive relationship observed in the rightmost plots in Figure 2, where the paired extreme curves appear somewhat similar. Second, by focusing on the extreme parts of paired samples of curves, our tool effectively quantifies risk during these critical events. The right plot of Figure 4 shows correlation coefficients computed from all curves, including non-extreme ones. These coefficients tend to underestimate the extreme risk, highlighting the necessity for tools that specifically describe extreme conditions.
@@ 7 Discussion
We introduced a coefficient designed to measure extremal dependence in paired samples of functions. This coefficient specifically focuses on extreme curves and computes their inner product to provide a quantitative assessment of the risk that paired extreme curves will occur simultaneously. The new coefficient is based on the theory of regular variation in $L^{2}$, providing a solid foundation for its derivation. Its estimator is shown to be consistent, a result supported by simulations. Deriving its asymptotic distribution might lead to a method for constructing confidence intervals. On the practical side, the estimator’s performance may depend on the method used to select an optimal $k$, as discussed in Section 5. While we considered two methods based on KS distances, that produce satisfactory results, future work could explore alternative methods and refine the selection of $k$.
Despite promising results, there are some limitation to our approach that suggest potential directions for future work. We treat functional observations as regularly varying square integrable random functions in $L^{2}$, which requires $\alpha>2$. While this condition is met for the financial and environmental data we work with, it might be desirable to find a correlation-like extremal dependence measure that requires merely $\alpha>0$. Possible extensions could involve the codifference or the covariation introduced in Kokoszka & Taqqu (1995). These measures of dependence are applicable to stable vectors with the index $\alpha<2$, and have been studied in econometrics and statistical physics contexts, see e.g., Kokoszka & Taqqu (1996) and Levy & Taqqu (2014) and Wolymańska et al. (2015). Exploring extreme value theory for functional data in this context might be useful, but one must note that those measures are not symmetric.
@@ Acknowledgement
We Thank Professor Hong Miao of the Department of Finance and Real Estate at Colorado State University for preprocessing the financial data used in Section 6.1. We thank Professor Joshua French of the Department of Mathematical and Statistical Sciences at the University of Colorado Denver for preprocessing the temperature data used in Section 6.2. Kim and Kokoszka were partially supported by the United States National Science Foundation.
@@ Appendix A Preliminary results
In this section, we put together preliminary results needed to prove Theorem 4.4. Some of these results are known in the literature, and none of them are particularly profound or difficult to prove. However, these results allow us to streamline the exposition of proofs of the main result. Recall that, c.f., (5), $\mathcal{A}_{r}=\{(x,y)\in{\mathbb{B}}_{0}^{2}:\|(x,y)\|_{{\mathbb{B}}^{2}}\geq r\}$, $r>0$, where $\|(x,y)\|_{{\mathbb{B}}^{2}}=\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}$.
@@
Let $M_{+}(0,\infty]$ be the space of Radon measures on $(0,\infty]$, and $\nu_{\alpha}(r,\infty]=r^{-\alpha}$. Also, let $\epsilon_{x}(A)=1$ if $x\in A$ and $\epsilon_{x}(A)=0$ if $x\notin A$. If $[X,Y]^{\top}$ is regularly varying in $L^{2}\times L^{2}$ according to Definition 3.1, then
(i) $R$ is a nonnegative random variable whose distribution has a regularly varying tail with index $-\alpha$,
@@
(iii) $R_{(k)}/b(n/k)\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}1$, in $[0,\infty)$,
@@
The following lemma is used to prove Lemmas A.6 and A.8.
@@
$\limsup_{n\to\infty}A_{n,j}^{(1)}\leq M^{2}\limsup_{n\to\infty}\gamma_{n}(K_{j}\setminus K^{\circ})\leq M^{2}\nu_{\alpha}(K_{j}\setminus K^{\circ})$
@@
[…]
Proof A.11.
@@
$V_{n,k}=\int_{1}^{\infty}r^{2}\hat{\gamma}_{n,k}(dr),$ $V=\int_{1}^{\infty}r^{2}\nu_{\alpha}(dr);$
$V_{n,k}^{(M)}=\int_{1}^{M}r^{2}\hat{\gamma}_{n,k}(dr),$ $V^{(M)}=\int_{1}^{M}r^{2}\nu_{\alpha}(dr).$
@@
[…]
@@
where $Z_{i}^{\prime}$ is an independent copy of $Z_{i}$ conditional on $Z_{j}=z_{j}$, $11\leq j\leq i-$ 1. If $b$ and $\hat{v}$ are finite, then for any $\varepsilon\geq 0$,
@@
Note that $\sigma_{n,k}$ is not observable since $b(\cdot)$ is unknown. However, $b(n/k)$ can be estimated by its consistent estimator $R_{(k)}$, and it can be shown that replacing $b(n/k)$ by $R_{(k)}$ ensures that the difference between $\sigma_{n,k}$ and $\hat{\sigma}_{n,k}$ is asymptotically negligible, which will be shown in Proposition B.3. Thus, the key argument for establishing the consistency is to show that $\sigma_{n,k}$ converges in probability to $\sigma_{XY}$, which is proven in the following proposition.
@@
If we apply this inequality to $-f({\bf Z}_{n})$, then we obtain the following ‘two-sided’ inequality
@@
$\leq\frac{\sqrt{n}}{k}\left\{E\left[\Delta_{1}^{2}\right]\right\}^{1/2}$
$=\frac{\sqrt{n}}{k}\left\{E\left[\left(\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle I_{R_{1}\geq b(n/k)}-E\left[\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle I_{R_{1}>b(n/k)}\right]\right)^{2}\right]\right\}^{1/2}$
$=\frac{\sqrt{n}}{k}\left\{\hbox{\rm var}\left[\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle I_{R_{1}\geq b(n/k)}\right]\right\}^{1/2}$
$\leq\frac{\sqrt{n}}{k}\left\{E\left[\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle^{2}I_{R_{1}\geq b(n/k)}\right]\right\}^{1/2}$
$\leq\frac{\sqrt{n}}{k}\left\{E\left[\left(\frac{R_{1}}{b(n/k)}\right)^{2}I_{R_{1}\geq b(n/k)}\right]\right\}^{1/2}.$

### v3

@@ 5 A simulation study
Assume that curves are observed on the regularly spaced grid $\{j/J,\ j\in\{1,\ldots,J\}\}$ on $[0,1]$, with each point assigned equal weight. The inner product and norm in $L^{2}([0,1])$ are then
@@
$g_{X}(x)=\frac{x}{\|x\|^{1-{\alpha_{X}}/\alpha}};\ \ \ \ g_{Y}(y)=\frac{y}{\|y\|^{1-{\alpha_{Y}}/\alpha}},\ \ \ \ x,y\in L^{2},$ (16)
where $\alpha$ is a desired tail index. Applying $g_{X}$ and $g_{Y}$ to $X$ and $Y$, respectively, ensures that $P(\|g_{X}(X)\|>\cdot)$ and $P(\|g_{Y}(Y)\|>\cdot)$ are regularly varying with the same index $\alpha$. Since this method adjusts only the scale of curves, their shapes are preserved.
Step 4. Given tail equivalent marginals, take the $k$ largest pairs of $X_{i},Y_{i}]^{\top}$, based on their norm, i.e., $R_{i}=\|X_{i}\|\vee\|Y_{i}\|$, and then compute $\hat{\rho}_{n,k}$ as in (14). By Lemma A.3 (i), if $X,Y]^{\top}$ are regularly varying in $L^{2}\times L^{2}$ with index $\alpha$, then $R=\|X\|\vee\|Y\|$ is regularly varying in $\mathbb{R}_{+}$ with the same index. Therefore, we choose $k$ that results in successful tail estimation for $R$ in finite samples. In the literature on tail estimation, various methods for selecting $k$ have been introduced, e.g., Hall & Welsh (1985)Hall (1990)Drees & Kaufmann (1998)Danielsson et al. (2001), just to name a few. We use the method of Danielsson et al. (2016), which is implemented using the function mindist of the R package tea.
We now outline the design of our simulation study. We generate functional observations in such a way that the theoretical value of $\rho_{XY}$ can be computed analytically, so that we can see how close $\hat{\rho}_{n,k}$ is to the true value. Suppose that $Z_{1}$ and $Z_{2}$ are i.i.d. random variables in $\mathbb{R}$} satisfying ${\rm pr}(|Z_{1}|>z)=z^{-\alpha}$ with equal chance of $Z_{1}$ being either negative or positive. Also, let $N_{1}$, $N_{2}$, and $N_{3}$ be i.i.d. normal random variables in $\mathbb{R}$} with mean 0 and variance 0.5. Consider $\{\phi_{j},j\geq 1\}$ defined by (3) and recall that it is an orthonormal basis in $L^{2}([0,1])$. These functions are simulated on a grid of 100 equally–spaced points on the unit interval $[0,1]$. We consider the following data generating processes, for $1\leq\rho\leq 1$,
$X(t)$ $=Z_{1}\phi_{1}(t)+N_{1}\phi_{2}(t)+N_{2}\phi_{3}(t);$ (17)
$Y(t)$ $=\rho Z_{1}\phi_{1}(t)+(1-\rho^{2})^{1/2}Z_{2}\phi_{2}(t)+N_{3}\phi_{3}(t).$
@@
Let $Z_{1},Z_{2}]^{\top}$ be a random vector in ${\mathbb{R}}^{2}$ consisting of i.i.d. components $Z_{j}$ whose magnitude is regularly varying with $\alpha$, $\alpha>2$, i.e., for some $c_{+}$, $c_{-}\geq 0$,
${\rm pr}(Z_{1}>z)\sim c_{+}z^{-\alpha},\quad\quad{\rm pr}(Z_{1}<-z)\sim c_{-}z^{-\alpha},$
where $f(z)\sim g(z)$ if and only if $\lim_{z\to\infty}f(z)/g(z)=1$. Also, let $\{\phi_{j},\ j\geq 1\}$ be a set of orthonormal elements in ${\mathbb{S}}_{L^{2}}$. Then, for $X$ and $Y$ in (17),
@@
We consider $\rho_{XY}\in\{0,\pm 0.1,\pm 0.2,\ldots,\pm 0.9,\pm 1\}$ and $\alpha\in\{3,4,5\}$, from which values of $\rho$ can be obtained by Lemma 5.1. For each $\rho$, we generate $X_{i},Y_{i}]^{\top}$, $1\leq i\leq n$, that are i.i.d. copies of $X,Y]^{\top}$, with sample sizes $n\in\{100,500,2000\}$. In each case, 1000 replications are generated. By Proposition 7.1 and Example 7.3 of Meinguet & Segers (2010), $X$ and $Y$ in (17) are regularly varying with the same index $\alpha$, so we proceed directly to Step 4 and compute $\hat{\rho}_{n,k}$. When choosing $k$, we also consider an alternative approach based on the Kolmogorov–Smirnov (KS) distance, as introduced by Clauset et al. (2009). The key difference from Danielsson et al. (2016) is that Clauset et al. (2009) computes the distance from tail distributions, rather than tail quantiles. The method is implemented using poweRlaw R package.
@@
We report the magnitude of empirical biases, measured by the absolute difference between the average and the theoretical value, along with standard errors computed as the sample standard deviations. Using the optimal $k$s selected by the method from Danielsson et al. (2016), the results are shown in Table 2 for $\alpha=3$. The results for $\alpha\in\{4,5\}$, as well as those using the method from Clauset et al. (2009), are provided in Section D.1 of Supplementary material, as they lead to similar conclusions. The conclusions are summarized as follows.
(i) The estimator is consistent as the biases decrease with increasing sample sizes, across nearly all values of $\rho_{XY}$ and for all $\alpha\in\{3,4,5\}$ considered. However, the finite performance using the optimal $k$s appears to depend on the tail index $\alpha$, with smaller $\alpha$ resulting in small biases.
(ii) The bias tends to increase in magnitude as $|\rho_{XY}|$ approaches 1. This could be due to the effect of the boundary, $\rho_{XY}\in\{-1,1\}$. These barriers cause the bias to increase, although it gets lower again at the boundary.
(iii) The standard errors are observed to be non-uniform across $\rho_{XY}$, they roughly behave like a quadratic function of $\rho_{XY}$ with its peak at $\pm.5$.
@@ 6.1 Extremal dependence of intraday returns on sector ETFs
Ticker Sector $\hat{\alpha}$ Ticker Sector $\hat{\alpha}$
XLY Consumer 3.8 XLV Health Care 3.9
Discretionary XLI Industrials 3.7
XLP Consumer Staples 2.6 XLB Materials 3.4
XLE Energy 4.2 XLK Technology 4.7
XLF Financials 4.0 XLU Utilities 3.8
@@
Figure 3 shows estimates of the pairwise extremal correlation coefficient across the nine ETF sectors. All pairs exhibit positive extremal correlations ($\hat{\rho}_{n,k}=0.43\sim 0.96$), and 44% of the pairs have strong extremal correlations above 0.7. We see that the CIDRs overall exhibit matching patterns of cumulative intraday returns on extreme market volatility days during the covid-19 market turbulence, where most sectors either drop or increase together. However, our coefficient reveals more subtle information as well. For example, extreme return curves of XLF (Financials) are exceptionally strongly correlated with extreme curves for XLV, XLB and XLK (Health Care, Materials, Technology), but moderately correlated with XLU (Utilities). We do not aim at an analysis of the stock market or the economy, but we observe that some findings are interesting. One might expect the financial sector (mostly banks) to be strongly affected by the technology sector (mostly large IT companies like Google or Microsoft) because such mega corporations dominate the U.S. stock market. The similarity of extreme return curves for XLF and XLK is shown in the leftmost panels of Fig. 2. In contrast, bank stocks would be expected to be less affected by utility companies, whose revenues are largely fixed. However, the strong correlation with the Health Care and Materials sectors is less clear. As another comparison, XLY (Consumer Discretionary) and XLU (Utilities) show a moderate extremal correlation of 0.43. Their extreme curves exhibit relatively dissimilar patterns, as seen in the middle panels in Fig. 2.
@@
We conclude by emphasizing that our tool offers a more precise quantification of intraday risk during extreme events. First, by analyzing curve shapes, it provides a better assessment of intraday risks. The left plot of Figure 4 displays the pairwise coefficients from single-valued closing returns. This plot reveals somewhat different information from Fig. 3. For instance, between XLY and XLE, the closing returns show a 0.3-0.3 correlation, indicating negative correlation during extreme days. However, this value does not accurately reflect the positive relationship observed in the rightmost plots in Fig. 2, where the paired extreme curves appear somewhat similar. Second, by focusing on the extreme parts of paired samples of curves, our tool effectively quantifies risk during these critical events. The right plot of Fig. 4 shows correlation coefficients computed from all curves, including non-extreme ones. These coefficients tend to underestimate the extreme risk, highlighting the necessity for tools that specifically describe extreme conditions.
@@ 7 Discussion
Despite promising results, there are some limitation to our approach that suggest potential directions for future work. We treat functional observations as regularly varying square integrable random functions in $L^{2}$, which requires $\alpha>2$. While this condition is met for the financial and environmental data we work with, it might be desirable to find a correlation-like extremal dependence measure that requires merely $\alpha>0$. Possible extensions could involve the codifference or the covariation introduced in Kokoszka & Taqqu (1995). These measures of dependence are applicable to stable vectors with the index $\alpha<2$, and have been studied in econometrics and statistical physics contexts, see e.g., Kokoszka & Taqqu (1996), Levy & Taqqu (2014), and Wolymańska et al. (2015). Exploring extreme value theory for functional data in this context might be useful, but one must keep in mind that those measures are not symmetric.
@@ Acknowledgement
We Thank Professor Hong Miao of the Department of Finance and Real Estate at Colorado State University for preprocessing the financial data used in Section 6.1. We thank Professor Joshua French of the Department of Mathematical and Statistical Sciences at the University of Colorado Denver for preprocessing the temperature data used in Section 6.2. We benefited from specific and detailed advice given by the reviewers. Kim and Kokoszka were partially supported by the United States National Science Foundation.
@@ Appendix A Preliminary results
In this section, we put together preliminary results needed to prove Theorem 4.4. Some of these results are known in the literature, and none of them are particularly profound or difficult to prove. However, these results allow us to streamline the exposition of proofs of the main result. Recall from (6) that $\mathcal{A}_{r}=\{(x,y)\in{\mathbb{B}}_{0}^{2}:\|(x,y)\|_{{\mathbb{B}}^{2}}\geq r\}$, $r>0$, where $\|(x,y)\|_{{\mathbb{B}}^{2}}=\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}$.
@@
Let $M_{+}(0,\infty]$ be the space of Radon measures on $(0,\infty]$, and $\nu_{\alpha}(r,\infty]=r^{-\alpha}$. Also, let $\epsilon_{x}(A)=1$ if $x\in A$ and $\epsilon_{x}(A)=0$ if $x\notin A$. If $X,Y]^{\top}$ is regularly varying in $L^{2}\times L^{2}$ according to Definition 3.1, then
(i) $R$ is a nonnegative random variable whose distribution has a regularly varying tail with index $\alpha$,
@@
(iii) $R_{(k)}/b(n/k)\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}1$, in $0,\infty)$,
@@
The following lemma is used to prove Lemmas A.8 and A.10.
@@
$\limsup_{n\to\infty}A_{n,j}^{(1)}\leq c\limsup_{n\to\infty}\gamma_{n}(K_{j}\setminus K^{\circ})\leq c\nu_{\alpha}(K_{j}\setminus K^{\circ})$
@@
Since $\tau$ is arbitrary, we get the claim.
The following lemma is used to prove Proposition 4.2.
Lemma A.6.
Under assumptions in Proposition 4.2, we have that
$\int_{{\mathcal{A}}_{1}}\left\langle x,y\right\rangle n{\rm pr}({b(n)}^{-1}X\in dx,{b(n)}^{-1}Y\in dy)\to\int_{{\mathcal{A}}_{1}}\left\langle x,y\right\rangle\mu(dx,dy).$
Proof A.7.
Since the function $(x,y)\mapsto\left\langle x,y\right\rangle I_{{\mathcal{A}}_{1}}$ is not bounded, we employ an approximation argument, similar to that in Lemma A.4. Let $\gamma_{n}=n{\rm pr}({b(n)}^{-1}X\in\cdot,{b(n)}^{-1}Y\in\cdot)$, and consider a sequence $\{{\mathcal{A}}_{j}\}_{j=1}^{\infty}$. By the triangle inequality, we have that
$\left|\int_{{\mathcal{A}}_{1}}\left\langle x,y\right\rangle\gamma_{n}(dx,dy)-\int_{{\mathcal{A}}_{1}}\left\langle x,y\right\rangle\mu(dx,dy)\right|$ $\leq\left|\int_{{\mathcal{A}}_{1}}\left\langle x,y\right\rangle\gamma_{n}(dx,dy)-\int_{{\mathcal{A}}_{1}\setminus{\mathcal{A}}_{j}}\left\langle x,y\right\rangle\gamma_{n}(dx,dy)\right|$
$\ \ +\left|\int_{{\mathcal{A}}_{1}\setminus{\mathcal{A}}_{j}}\left\langle x,y\right\rangle\gamma_{n}(dx,dy)-\int_{{\mathcal{A}}_{1}\setminus{\mathcal{A}}_{j}}\left\langle x,y\right\rangle\mu(dx,dy)\right|$
$\ \ +\left|\int_{{\mathcal{A}}_{1}\setminus{\mathcal{A}}_{j}}\left\langle x,y\right\rangle\mu(dx,dy)-\int_{{\mathcal{A}}_{1}}\left\langle x,y\right\rangle\mu(dx,dy)\right|$
$=:B_{n,j}^{(1)}+B_{n,j}^{(2)}+B_{j}^{(3)}.$
Fix $\tau>0$. It follows by (11) that there is $j^{\star}$ such that for $j\geq j^{\star}$,
$B_{j}^{(3)}\leq\int_{{\mathcal{A}}_{j}}\left\langle x,y\right\rangle\mu(dx,dy)=\int_{r\geq j}r^{2}\alpha r^{-\alpha-1}dr\int_{{\mathbb{S}}_{L^{2}\times L^{2}}}\left\langle\theta_{X},\theta_{Y}\right\rangle\Gamma(d\theta_{X},d\theta_{Y})\leq\frac{\alpha}{\alpha-2}j^{-\alpha+2}<\tau/2.$
Similarly, it follows by change of variable that
$B_{n,j}^{(1)}$ $\leq\int_{{\mathcal{A}}_{j}}\left\langle x,y\right\rangle\gamma_{n}(dx,dy)$
$=\int_{r\geq j}\int_{{\mathbb{S}}_{L^{2}\times L^{2}}}r^{2}\left\langle\theta_{X},\theta_{Y}\right\rangle n{\rm pr}\left(\frac{\|(X,Y)\|}{b(n)}\in dr,\frac{(X,Y)}{\|(X,Y)\|}\in(d\theta_{X},d\theta_{Y})\right)$
$\leq\int_{r\geq j}r^{2}\ n{\rm pr}\left(\frac{\|(X,Y)\|}{b(n)}\in dr\right).$
Then, by Lemmas A.3 (i) and A.4, we have that for each $j\geq j^{\star}$
$\limsup_{n\to\infty}B_{n,j}^{(1)}\leq\int_{r\geq j}r^{2}\nu_{\alpha}(dr)=\frac{\alpha}{\alpha-2}j^{-\alpha+2}<\tau/2.$
Thus, it follows from Lemma A.1 and Lemma A.1 of Meinguet & Segers (2010) that
$\limsup_{n\to\infty}\left|\int_{{\mathcal{A}}_{1}}\left\langle x,y\right\rangle\gamma_{n}(dx,dy)-\int_{{\mathcal{A}}_{1}}\left\langle x,y\right\rangle\mu(dx,dy)\right|\leq\tau+\limsup_{n\to\infty}B_{n,j^{\star}}^{(2)}=\tau.$
@@
$\frac{1}{k}\sum_{i=1}^{n}\left(\frac{R_{i}}{R_{(k)}}\right)^{2}I_{R_{i}\geq R_{(k)}}$ $\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}\frac{\alpha}{\alpha-2};$ (18)
$\frac{1}{k}\sum_{i=1}^{n}\left(\frac{R_{i}}{b(n/k)}\right)^{2}I_{R_{i}\geq b(n/k)}$ $\stackrel{{\scriptstyle\mbox{$\scriptstyle P$}}}{{\rightarrow}}\frac{\alpha}{\alpha-2}.$ (19)
Proof A.13.
@@
$V_{n,k}=\int_{1}^{\infty}r^{2}\hat{\gamma}_{n,k}(dr),\ \ \$$V=\int_{1}^{\infty}r^{2}\nu_{\alpha}(dr);$
$V_{n,k}^{(M)}=\int_{1}^{M}r^{2}\hat{\gamma}_{n,k}(dr),\ \ \$$V^{(M)}=\int_{1}^{M}r^{2}\nu_{\alpha}(dr).$
@@
$\int_{M}^{\infty}r^{2}\nu_{\alpha}(dr)=\int_{M}^{\infty}r^{2}\alpha r^{-\alpha-1}dr=\frac{\alpha}{\alpha-2}M^{2-\alpha}\to 0,\ \ \ \ {\rm as}\ M\to\infty.$
@@
where $Z_{i}^{\prime}$ is an independent copy of $Z_{i}$ conditional on $Z_{j}=z_{j}$, $1\leq j\leq i-1$. If $b$ and $\hat{v}$ are finite, then for any $\varepsilon\geq 0$,
@@
We emphasize that $\sigma_{n,k}$ is not observable since $b(\cdot)$ is unknown. However, $b(n/k)$ can be estimated by its consistent estimator $R_{(k)}$, and it can be shown that replacing $b(n/k)$ by $R_{(k)}$ ensures that the difference between $\sigma_{n,k}$ and $\hat{\sigma}_{n,k}$ is asymptotically negligible, which will be shown in Proposition B.3. Thus, the key argument for establishing the consistency is to show that $\sigma_{n,k}$ converges in probability to $\sigma_{XY}$, which is proven in the following proposition.
@@
If we apply this inequality to $f({\bf Z}_{n})$, then we obtain the following ‘two-sided’ inequality
@@
$\leq\frac{n^{1/2}}{k}\left\{E\left[\Delta_{1}^{2}\right]\right\}^{1/2}$
$=\frac{n^{1/2}}{k}\left\{E\left[\left(\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle I_{R_{1}\geq b(n/k)}-E\left[\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle I_{R_{1}>b(n/k)}\right]\right)^{2}\right]\right\}^{1/2}$
$=\frac{n^{1/2}}{k}\left\{\hbox{\rm var}\left[\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle I_{R_{1}\geq b(n/k)}\right]\right\}^{1/2}$
$\leq\frac{n^{1/2}}{k}\left\{E\left[\left\langle\frac{X_{1}}{b(n/k)},\frac{Y_{1}}{b(n/k)}\right\rangle^{2}I_{R_{1}\geq b(n/k)}\right]\right\}^{1/2}$
$\leq\frac{n^{1/2}}{k}\left\{E\left[\left(\frac{R_{1}}{b(n/k)}\right)^{2}I_{R_{1}\geq b(n/k)}\right]\right\}^{1/2}.$